# MOMEMTO: Patch-based Memory Gate Model in Time Series Foundation Model

### v1

@@ Abstract
Recently reconstruction-based deep models have been widely used for time series anomaly detection, but as their capacity and representation capability increase, these models tend to over-generalize, often reconstructing unseen anomalies accurately. Prior works have attempted to mitigate this by incorporating a memory architecture that stores prototypes of normal patterns. Nevertheless, these approaches suffer from high training costs and have yet to be effectively integrated with time series foundation models (TFMs). To address these challenges, we propose MOMEMTO, a TFM for anomaly detection, enhanced with a patch-based memory module to mitigate over-generalization. The memory module is designed to capture representative normal patterns from multiple domains and enables a single model to be jointly fine-tuned across multiple datasets through a multi-domain training strategy. MOMEMTO initializes memory items with latent representations from a pre-trained encoder, organizes them into patch-level units, and updates them via an attention mechanism. We evaluate our method using 23 univariate benchmark datasets. Experimental results demonstrate that MOMEMTO, as a single model, achieves higher scores on AUC and VUS metrics compared to baseline methods, and further enhances the performance of its backbone TFM, particularly in few-shot learning scenarios.
@@ 1 Introduction
Recent research has increasingly focused on TFMs, which are pre-trained on large-scale time series data from a variety of sources. TFMs can be adapted to new tasks with minimal additional training through approaches such as multi-task learning, few-shot learning, and prompting. However, applying TFMs to anomaly detection presents two main challenges. First, most TFMs, such as MOIRAI (Woo et al., 2024), TimesFM (Das et al., 2024), Chronos (Ansari et al., 2024), Time-MoE (Shi et al., 2024), and Sundial (Liu et al., 2025) are primarily designed for time series forecasting. They often employ decoder-only architectures that limit their suitability for detection tasks. Second, TFMs used for anomaly detection typically adopt reconstruction-based approaches, leaving them vulnerable to over-generalization.
To address these challenges, we adapt the Gate memory module of MEMTO (Song et al., 2023), a reconstruction-based deep model for time series anomaly detection. This module stores memory items that represent the prototypical features of normal patterns. When anomalies are reconstructed using memory items containing normal features, the resulting outputs resemble normal samples, thus alleviating over-generalization. However, MEMTO is highly sensitive to memory initialization, requiring a pre-training stage for the encoder to learn sufficiently informative representations. In addition, MEMTO lacks a patch-based design and follows a one-model-per-dataset framework. These characteristics lead to high training costs when handling multiple datasets.
@@
Building upon previous methods, we introduce MOMEMTO, a new reconstruction-based TFM for time series anomaly detection. The model integrates the pre-trained encoder from MOMENT with the patch-based memory module. Unlike MEMTO, our model benefits from MOMENT’s encoder, which already possesses sufficient representational capacity. This provides well-initialized features for the memory module and alleviates MEMTO’s sensitivity to memory initialization. The patch-based memory module stores encoder outputs that are organized at the patch level. Whereas the original memory items stored prototypical features of normal patterns restricted to a single dataset, the patch-based memory items are designed to capture representative normal patterns from multiple domains. This design alleviates over-generalization while enabling the model to learn effectively in diverse domains. Specifically, multi-domain training refers to jointly fine-tuning a single model across datasets from diverse application areas, rather than training a separate model for each dataset. This strategy facilitates knowledge sharing across domains and allows more effective use of limited data, while also reducing computational cost in both training time and memory usage. Under the multi-domain training strategy, a single MOMEMTO model achieves strong results on 23 univariate benchmark datasets compared to baseline methods. In few-shot learning scenarios, it demonstrates significant improvements over its backbone TFM. Our contributions can be summarized as follows:
@@
MOMEMTO is the first TFM specialized in time series anomaly detection, supporting multi-domain training and leveraging the patch-based memory module to mitigate over-generalization.
@@ 2 Related Work
MOMEMTO extends MEMTO by refining its memory mechanism to explicitly capture patch-level patterns. Our patch-based memory module is designed based on patching and channel-independent strategies, allowing selective updates of memory items rather than updating all items as in MEMTO. This design enables the memory module to learn more informative features of normal patterns while also improving the computational efficiency of memory updates compared to the original memory module (Nie et al., 2023; Song et al., 2023).
@@ 3.1.2 Patch-based Memory Module
To alleviate over-generalization in reconstruction-based models, we introduce the patch-based memory module as a redesigned variant of MEMTO’s original memory module. It naturally aligns with the encoder’s patch representations. The module consists of memory items , where each memory item stores a prototypical normal pattern at the patch level. The module operates in a data-driven manner through four stages: memory initialization, memory alignment, memory update, and query update, which iteratively align memory items with representative normal patterns.
@@
After initializing the memory items, we first identify the observed patches in the encoder output using the input mask and retrieve the corresponding subset of representations. To ensure consistent interaction, each memory item is aligned with the same subset of observed patches, and the queries are likewise L2-normalized, denoted as .
@@
We then compute the similarity between the normalized queries and the corresponding memory slices using dot products. Instead of updating every memory item, we select the top- memory items with the highest similarity scores. These selected memory items are updated through query-based patch-level attention, where denotes the memory item with the -th highest similarity to . The resulting updates are proportionally applied to the selected memory items in a data-driven way, with the remaining update procedure following MEMTO’s standard strategy (details in Appendix B.4).
@@
where denotes the similarity score between and , computed in the previous stage. The resulting is then used jointly with the original as input to the decoder.
@@ 4 Experiments
We evaluate our model on the TSB-AD-U benchmark (Liu & Paparrizos, 2024), consisting of 870 univariate time series across 23 datasets. The benchmark is constructed through an organized curation process: both univariate and multivariate datasets are collected, with multivariate data converted into univariate channels. Channels unrelated to anomaly labels, such as categorical, binary, or noisy channels, are discarded. Mislabeled or low-quality series are removed through algorithmic tests and human review, and balanced sampling is applied to ensure fairness. The collection also includes sequences from public benchmarks such as MSL (Mars Science Laboratory rover), SMD (Server Machine Dataset), and the UCR Anomaly Archive (Wu & Keogh, 2022). Further dataset details are provided in Appendix A.
@@
We generate subsequences by applying a non-overlapping window with a length of 512 to obtain fixed-length inputs for each time series. We evaluate anomaly detection performance using threshold-independent metrics, such as the Area Under Curve (AUC) and Volume Under the Surface (VUS) (Paparrizos et al., 2022). More detailed information on hyperparameter settings can be found in Appendix B.
@@ 4.1 Main Results
In this experiment, we evaluate the performance of MOMEMTO on time series anomaly detection by comparing it with 13 baseline methods. The baselines consist of three categories: classic approaches (OCSVM (Schölkopf et al., 1999), LOF (Breunig et al., 2000), Isolation Forest (Liu et al., 2008)), Deep learning models (LSTM-AD (Malhotra et al., 2015), Donut (Xu et al., 2018), OmniAnomaly (Su et al., 2019), TranAD (Tuli et al., 2022), Anomaly Transformer (Xu et al., 2022), MEMTO, TimesNet (Wu et al., 2023)), and pre-trained TFMs (TimesFM, Chronos, MOMENT).
@@ 4.2 Model Analysis
We visualize in Figure 4 how memory items are accessed for subsequences from different domains. Each heatmap displays the accumulated similarity between the true domain of the input and the selected memory domains, normalized such that each row sums to one. The left illustrates updates during training, and the right shows references by new inputs at test time. With the data-driven strategy, no domain labels are provided during training or testing. Updates tend to concentrate on a subset of memory items, rather than being evenly distributed across domains, and these items are consistently referenced at test time. To validate the effectiveness of our strategy, we compare it with two alternative strategies: (i) freezing the memory items without updates and (ii) updating only the memory item initially assigned to the input’s domain (own-domain update). Both baselines use the memory item corresponding to the input’s domain during training and testing. Table 3 summarizes the performance comparison. Even though the initial allocation of memory items to domains does not remain aligned during training, our data-driven strategy still achieves higher performance while utilizing fewer memory items than alternative strategies. This indicates that the memory module has learned a stable set of domain-general features, which are effectively reused across unseen inputs.
@@
Time (sec) Model size (GB) Training Context switching Total training Inference MOMEMTO 82.22 5.17 87.39 54.61 1.30 w/o PMM 75.83 4.59 80.42 54.23 1.29 w/o multi-domain training 135.77 1311.51 1447.28 54.54 455.74 w/o PMM & multi-domain training 108.96 1294.26 1403.22 54.29 451.64
@@ Appendix A TSB-AD-U Benchmark
In this section, we provide details of the TSB-AD-U benchmark (Liu & Paparrizos, 2024) used in our experiments. The benchmark comprises 23 univariate time series datasets with a total of 870 time series, collected from diverse domains. Table 5 summarizes the dataset composition, including the number of series, average sequence length, anomaly counts, and anomaly ratios. The benchmark is divided into two partitions: the Eval set, which contains 350 time series and is designated for evaluation, and the Tuning set, which contains 48 time series and is used for hyperparameter optimization. Table 6 summarizes the statistics of the Eval and Tuning sets.
@@ C.3 Number of Referenced Memory Items
C.4 Anomaly Score Distributions

### v3

@@ Abstract
Recently reconstruction-based deep models have been widely used for time series anomaly detection, but as their capacity and generalization capability increase, these models tend to over-generalize, often reconstructing unseen anomalies accurately. Prior works have attempted to mitigate this by incorporating a memory architecture that stores prototypes of normal patterns. Nevertheless, these approaches suffer from high training costs and have yet to be effectively integrated with time series foundation models (TFMs). To address these challenges, we propose MOMEMTO, an improved variant of TFM for anomaly detection, enhanced with a patch-based memory module to mitigate over-generalization. The memory module is designed to capture representative normal patterns from multiple domains and enables a single model to be jointly fine-tuned across multiple datasets through a multi-domain training strategy. MOMEMTO initializes memory items with latent representations from a pre-trained encoder, organizes them into patch-level units, and updates them via an attention mechanism. We evaluate our method using 23 univariate benchmark datasets. Experimental results demonstrate that MOMEMTO, as a single model, achieves higher scores on AUC and VUS metrics compared to baseline methods, and further enhances the performance of its backbone TFM, particularly in few-shot learning scenarios.
@@ 1 Introduction
Recent research has increasingly focused on TFMs, which are pre-trained on large-scale time series data from a variety of sources. Their strong generalization capability enables TFMs to achieve robust few-shot and zero-shot performance on new tasks. However, applying TFMs to anomaly detection presents two main challenges. First, most TFMs, such as MOIRAI (Woo et al., 2024), TimesFM (Das et al., 2024), Chronos (Ansari et al., 2024), Time-MoE (Shi et al., 2024), and Sundial (Liu et al., 2025) are primarily designed for time series forecasting. They often employ decoder-only architectures that limit their suitability for detection tasks. Second, the generalization capability of TFMs leaves them vulnerable to over-generalization under reconstruction approaches, paradoxically hindering their effectiveness as a unified model across diverse datasets.
To address these challenges, we adapt the Gate memory module of MEMTO (Song et al., 2023), a reconstruction-based deep model for time series anomaly detection. This module stores memory items that represent the prototypical features of normal patterns. When anomalies are reconstructed using these normal memory items, the resulting outputs resemble normal samples, thus alleviating over-generalization. However, MEMTO is highly sensitive to memory initialization, requiring encoder pre-training to obtain informative and stable representations. Given MEMTO’s one-model-per-dataset design, this sensitivity leads to high training cost when handling multiple datasets. It also uses a point-level memory rather than a patch-based structure, which reduces its effectiveness in detecting interval or periodic anomalies (Shen, 2025).
@@
Building upon previous methods, we introduce MOMEMTO, an improved variant of MOMENT adapted for reconstruction-based time series anomaly detection. The model integrates the pre-trained encoder from MOMENT with the patch-based memory module. Unlike MEMTO, our model benefits from MOMENT’s encoder, which already possesses sufficient representational capacity. This provides well-initialized features for the memory module and alleviates MEMTO’s sensitivity to memory initialization. The patch-based memory module stores encoder outputs that are organized at the patch level. Whereas the original memory items stored prototypical features of normal patterns restricted to a single dataset, the patch-based memory items are designed to capture representative normal patterns from multiple domains. This design alleviates over-generalization while enabling the model to learn effectively in diverse domains. Specifically, multi-domain training refers to jointly fine-tuning a single model across datasets from diverse application areas, rather than training a separate model for each dataset. This strategy facilitates knowledge sharing across domains and allows more effective use of limited data, while also reducing computational cost in both training time and memory usage. Under the multi-domain training strategy, a single MOMEMTO model achieves strong results on 23 univariate benchmark datasets compared to baseline methods. In few-shot learning scenarios, it demonstrates significant improvements over its backbone TFM. Our contributions can be summarized as follows:
@@
MOMEMTO extends MOMENT into a TFM specialized for time series anomaly detection, supporting multi-domain training and leveraging the patch-based memory module to mitigate over-generalization.
@@ 2 Related Work
MOMEMTO extends MOMENT by introducing a novel patch-based memory mechanism to explicitly capture patch-level patterns. Our patch-based memory module is designed based on patching and channel-independent strategies, preserving patch-level information, while allowing selective updates of memory items rather than updating all items. This design enables the memory module to learn more stable features of normal patterns while also improving the computational efficiency of memory updates compared to the prior memory methods (Nie et al., 2023; Song et al., 2023; Shen, 2025).
@@ 3.1.2 Patch-based Memory Module
To alleviate over-generalization in reconstruction-based models, we introduce the patch-based memory module as a redesigned variant of MEMTO’s original memory module. It naturally aligns with the encoder’s patch representations. The module consists of memory items , where each memory item stores a prototypical normal pattern at the patch level. The module operates in a data-driven manner through five stages: memory initialization, memory alignment, memory update, query update, and query alignment. Through these stages, the memory items and queries are progressively refined and aligned with representative normal patterns.
@@
After initializing the memory items, we first identify the observed patches in the encoder output using the input mask and retrieve the corresponding subset of representations. To ensure consistent interaction, each memory item is aligned with the same subset of observed patches, and the queries are likewise L2-normalized.
@@
We then compute the similarity between the normalized queries and the corresponding memory slices using dot products. Instead of updating every memory item, we select the top- memory items with the highest similarity scores. These selected memory items are updated through query-based patch-level attention, where denotes the memory item with the -th highest similarity to . The resulting updates are proportionally applied to the selected memory items in a data-driven way:
where are learnable projection matrices, and denotes the element-wise sigmoid (details in Appendix B.4).
@@
where denotes the similarity score between and , computed in the previous stage.
Query alignment
The resulting is then combined with the original in a patch-aligned manner, where each patch-level memory item contributes only to the corresponding patch position. This combined representation is then used as the input to the shared decoder. This ordered memory structure preserves the positional structure of the sequence and captures the underlying patterns in each patch.
@@ 4 Experiments
We evaluate our model on the TSB-AD-U benchmark (Liu and Paparrizos, 2024), consisting of 870 univariate time series across 23 datasets. The benchmark is constructed through an organized curation process: both univariate and multivariate datasets are collected, with multivariate data converted into univariate channels. Channels unrelated to anomaly labels, such as categorical, binary, or noisy channels, are discarded. Mislabeled or low-quality series are removed through algorithmic tests and human review, and balanced sampling is applied to ensure fairness. The collection also includes sequences from public benchmarks such as MSL (Mars Science Laboratory rover), SMD (Server Machine Dataset), and the UCR Anomaly Archive (Wu and Keogh, 2022). Further dataset details are provided in Appendix A.
@@
We generate subsequences by applying a non-overlapping window with a length of 512 to obtain fixed-length inputs for each time series. We evaluate anomaly detection performance using threshold-independent metrics, such as the Area Under the Curve (AUC) and Volume Under the Surface (VUS) (Paparrizos et al., 2022). More detailed information on hyperparameter settings can be found in Appendix B.
@@ 4.1 Main Results
In this experiment, we evaluate the performance of MOMEMTO on time series anomaly detection by comparing it with 15 baseline methods. The baselines consist of three categories: classical approaches (OCSVM (Schölkopf et al., 1999), LOF (Breunig et al., 2000), Isolation Forest (Liu et al., 2008)), Deep learning models (LSTM-AD (Malhotra et al., 2015), Donut (Xu et al., 2018), OmniAnomaly (Su et al., 2019), TranAD (Tuli et al., 2022), Anomaly Transformer (Xu et al., 2022), MEMTO, TimesNet (Wu et al., 2023), KAN-AD (Zhou et al., 2025)), and pre-trained TFMs (TimesFM, Chronos, MOMENT, DADA (Shentu et al., 2025)).
@@ 4.2 Model Analysis
We visualize in Figure 4 how memory items are accessed for subsequences from different domains. Each heatmap displays the accumulated similarity between the true domain of the input and the selected memory domains, normalized such that each row sums to one. The left illustrates updates during the entire training, and the right shows references by new inputs at test time. With the data-driven strategy, no domain labels are provided during training or testing. Updates tend to concentrate on a subset of memory items, rather than being evenly distributed across domains, and these items are consistently referenced at test time. To validate the effectiveness of our strategy, we compare it with two alternative strategies: (i) freezing the memory items without updates and (ii) updating only the memory item initially assigned to the input’s domain (own-domain update). Both baselines use the memory item corresponding to the input’s domain during training and testing. Table 3 summarizes the performance comparison. Even though the initial allocation of memory items to domains does not remain aligned during training, our data-driven strategy still achieves higher performance while utilizing fewer memory items than alternative strategies. This indicates that the memory module has learned a stable set of domain-general features, which are effectively reused across unseen inputs.
@@
Time (s) # Params (B) CKPT (GB) Training Context switching Total training Inference MOMEMTO 82.22 5.17 87.39 54.61 0.388 1.30 w/o PMM 75.83 4.59 80.42 54.23 0.385 1.29 w/o multi-domain training 135.77 1311.51 1447.28 54.54 135.80 455.74 w/o PMM & multi-domain training 108.96 1294.26 1403.22 54.29 134.75 451.64
@@ Appendix A TSB-AD-U Benchmark
In this section, we provide details of the TSB-AD-U benchmark (Liu and Paparrizos, 2024) used in our experiments. The benchmark comprises 23 univariate time series datasets with a total of 870 time series, collected from diverse domains. Table 5 summarizes the dataset composition, including the number of series, average sequence length, anomaly counts, and anomaly ratios. The benchmark is divided into two partitions: the Eval set, which contains 350 time series and is designated for evaluation, and the Tuning set, which contains 48 time series and is used for hyperparameter optimization. Table 6 summarizes the statistics of the Eval and Tuning sets.
@@ C.3 Number of Referenced Memory Items
C.4 Number of Memory Items
To evaluate how the number of memory items affects MOMEMTO, we conduct a sensitivity analysis by varying the number of memory items from 64 down to 1. In this experiment, each memory item is initialized using K-means centroid values.
| 64 | 32 | 16 | 8 | 4 | 2 | 1 | |
| AUC-PR | 36.11 | 36.05 | 36.26 | 36.07 | 35.97 | 35.69 | 35.81 |
| AUC-ROC | 74.59 | 74.22 | 74.18 | 74.19 | 74.30 | 73.25 | 73.50 |
Figure 6 visualizes the row-normalized scaled domain-to-memory similarities at the beginning and end of training. Across all configurations, the model consistently relies on a small subset of memory items, regardless of the total number of memory items.
| 0.1 | 0.2 | 0.3 | 0.5 | 0.7 | 0.9 | |
| 32 | 31.02 | 32.84 | 33.40 | 35.04 | 35.23 | 35.63 |
| 16 | 30.98 | 32.83 | 33.46 | 35.01 | 35.36 | 35.55 |
| 8 | 30.95 | 32.75 | 33.48 | 34.81 | 35.21 | 35.16 |
| 4 | 30.93 | 32.42 | 33.52 | 34.69 | 35.42 | 35.32 |
| 2 | 30.85 | 32.80 | 33.52 | 35.08 | 35.67 | 35.65 |
| 1 | 30.93 | 32.32 | 33.50 | 34.84 | 35.60 | 35.51 |
These results indicate that MOMEMTO exhibits stable performance with respect to the number of memory items.
C.5 Anomaly Score Distributions